OpenStudy (anonymous):

Ok, so the slope of the coordinates is (2,6) and (4,8). The slope is; M= 8-6 ---- 2/2 which equals 1. 4-2 I need help with; Take the points you have selected and write the equation of the line using the point-slope formula.

OpenStudy (anonymous):

too easy that is why you are confused

OpenStudy (anonymous):

you can write either \[y-6=x-2\] or use the other point because the slope is 1 you do not have to write \(y-y_1=1(x-x_1)\)

OpenStudy (anonymous):

Oh damn, I feel silly.

OpenStudy (anonymous):

lol i thought you might

OpenStudy (anonymous):

Could you show me how to show the work?

OpenStudy (anonymous):

Besides one other person, you have been the most reliable source lol.

OpenStudy (anonymous):

what work?

OpenStudy (anonymous):

According to the assignment, I need to show work.

OpenStudy (anonymous):

\[y-6=x-2\] or \[y-8=x-4\] same thing

OpenStudy (anonymous):

you did all the work yourself in finding the slope the rest is just to write the answer

OpenStudy (anonymous):

both equations are \(y=x+4\) if you solve for \(y\) but you were not asked to solve for \(y\) you were only asked for "point-slope" form, either is correct

OpenStudy (anonymous):

Ohhhh, ok thank you. I have a high C in this class so I have been working very hard to bring it up, so that is why I am being so bleh about this stuff lol.

OpenStudy (anonymous):

you can do this for sure

OpenStudy (anonymous):

Ok next I need to take my points (2,6) and (4,8) and do this; Take the points you have selected and write the equation of the line using the slope-intercept formula

OpenStudy (anonymous):

i thought you just did this one the slope is 1 the equation is either \(y-4=x-2\) or \(y-8=x-4\)

OpenStudy (anonymous):

I think I did, but this is the third question, this is just weird.

OpenStudy (anonymous):

looks the same to me

OpenStudy (anonymous):

Yeah, I know, hold on, let me type this out and see what is next.

OpenStudy (anonymous):

The last one was just a write out the steps, thank you for your help.